# Sums related to Euler’s totient function

### v1

@@ Abstract.
Euler’s totient function, prime numbers.
@@ 1. Introduction
Let $\varphi$ denote Euler’s totient function and $\log_{2}s=\max(\ln\ln s,1)$. Our main result is the following.
@@
Let $g$ be the positive multiplicative function such that $g(p)\geq c_{0}>$ for any prime $p$ and
$L=\sum_{p}\frac{1}{pg(p)}<\infty.$
@@
$\omega(n)\leq K\frac{N}{g(n)}$
for any $n\in\mathcal{D}$, where $K>0$ is a constant. Then
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)N$ (1.1)
for any $s\in\mathbb{N}$. Moreover,
#$\#\big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N.$ (1.2)
Here $C$, $c_{1}$, and $c_{2}$ are positive constants depending only on $K$, $L$, $c_{0}$, and $\alpha$.
From Theorem 1.1 we obtain the following results.
@@
$\sum_{p\leq x}\Big{(}\frac{p-1}{\varphi(p-1)}\Big{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)\pi(x)$
@@
#$\#\big{\{}p\leq x:\frac{p-1}{\varphi(p-1)}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$
Corollary 1.2.
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)x$ (1.3)
and
@@ 2. Notation
The symbol $b|a$ means that $b$ divides $a$. For fixed $a$ the sum $\sum_{b|a}$ and the product $\prod_{b|a}$ should be interpreted as being over all positive divisors of $a$. If $x$ is a real number, then $\lfloor x\rfloor$ denotes its integral part.
3. Proof of Theorem 1.1 and Corollaries 1.1 - LABEL:C3
@@
It is easy to see that $f(n)$ is a multiplicative function supported on $\mathcal{D}$ and
@@
Since $\omega(n)\leq KN/g(n)$ for any $n\in\mathcal{D}$, we have
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$ $\leq KN\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n\in\mathcal{D}}\frac{f(n)}{g(n)}=KN\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{f(p)}{g(p)}\Big{)}$
$\leq KN\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$ (3.1)
@@
$\Big{(}1+\frac{1}{p}\Big{)}^{s}-1=s(1+\xi)^{s-1}\frac{1}{p}<s\Big{(}1+\frac{1}{s}\Big{)}^{s-1}\frac{1}{p}<\frac{es}{p}.$
@@
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\textup{exp}(s\log_{2}s+Cs)N,$
where $C$ is a positive constant depending only on $K$, $L$, $c_{0}$, and $\alpha$. The inequality (1.1) is proved.
@@
$\textup{exp}(s\log_{2}s+Cs)N$ $\geq\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\geq\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(s\log_{2}s+Cs-s\ln t)N.$
@@
$s=\lfloor\textup{exp}(te^{-(C+1)})\rfloor+1.$
It is easy to see that
$s\log_{2}s+Cs-s\ln t=-\textup{exp}(te^{-(C+1)})+O(\ln t)\leq-\textup{exp}(te^{-(C+1)}/2),$
if $t\geq t_{0}$ (here $t_{0}=t_{0}(K,L,c_{0},\alpha)$ is a positive constant depending only on $K$, $L$, $c_{0}$, and $\alpha$). We set
@@
Then $c_{1}$ and $c_{2}$ are positive constants depending only on $K$, $L$, $c_{0}$, and $\alpha$. If $t\leq t_{0}$, then
@@
We set $a_{n}=p_{n}-1$, where $p_{n}$ is the $n$th prime. It is clear that $N=\pi(x)$. We take $\alpha=1/2$, $M=x$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is large constant. If $n\in\mathcal{D}$, then
By the Brun - Titchmarsh inequality we have
$\omega(n)=\#\{p\leq x:p\equiv 1\text{ {(}mod $n${)}}\}\leq\frac{3x}{\varphi(n)\ln(2x/n)}\leq K\,\frac{\pi(x)}{\varphi(n)},$
where $K$ is an absolute positive constant. Since $\varphi(n)$ is a positive multiplicative function, $\varphi(p)=p-1\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$), and
$L=\sum_{p}\frac{1}{p\varphi(p)}=\sum_{p}\frac{1}{p(p-1)}<\infty,$
we can apply Theorem 1.1. Since $K$, $L$, $c_{0}$, and $\alpha$ are absolute positive constants, the constants $C$, $c_{1}$, and $c_{2}$ are absolute and positive. Corollary 1.1 is proved. ∎
Proof of Corollary 1.2..
We set $a_{n}=f(n)$. It is clear that $N=\lfloor x\rfloor$. There exists a positive constant $\gamma$ depending only on $f$ such that $f(n)\leq\gamma n^{d}$ for any positive integer $n$. We take $M=\gamma x^{d}$ and $\alpha=1/2$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is large constant depending only on $f$. We have
$y=(\ln M)^{\alpha}=(d\ln x+\ln\gamma)^{1/2}\leq(\ln x)^{3/4},$
@@
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(2(\ln x)^{3/4})\leq\sqrt{x}<\lfloor x\rfloor=N.$
@@
$\leq(2c)dn^{1-1/d}\,\frac{N}{n}=K\,\frac{N}{n^{1/d}},$
where $K=2cd$. The function $g(n)=n^{1/d}$ is positive and multiplicative. Also, $g(p)=p^{1/d}\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$) and

### v2

@@ Abstract.
Euler’s totient function, multiplicative functions, prime numbers.
@@ 1. Introduction
Let $\varphi$ denote Euler’s totient function. We prove
@@
$\omega(n)\leq\frac{K}{g(n)}$
for any $n\in\mathcal{D}$, where $K>0$ is a constant (depending on $N$) and $g(n)$ is a multiplicative function. Then
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
for any $s\in\mathbb{N}$. Here $c>0$ is an absolute constant.
Theorem 1.2.
Suppose that the assumptions of Theorem 1.1 hold, $K=\gamma N$, $g(p)\geq c_{0}>0$ for any prime $p$, and
$L=\sum_{p}\frac{1}{pg(p)}<\infty.$
Then there are positive constants $C$, $c_{1}$, and $c_{2}$ depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$ such that
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)N$ (1.1)
for any $s\in\mathbb{N}$, and
#$\#\big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))N$ (1.2)
for any $t>0$.
From Theorems 1.1 and 1.2 we obtain the following results.
Corollary 1.1.
$\sum_{n\leq x}\Big{(}\frac{f(n)}{\varphi(f(n))}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)x$ (1.3)
and
Corollary 1.1 extends Corollary 1.1 in [3] which showed the inequality (1.3) but with an upper bound $\textup{exp}(s\ln s+Cs)x$ and without the inequality (1.4).
Let =$\mathcal{L}=\{L_{1},\ldots,L_{k}\}$ be a set of $k$ linear functions with integer coefficients
$L_{i}(n)=a_{i}n+b_{i},\qquad i=1,\ldots,k.$
For $L(n)=an+b$, $a,b\in\mathbb{Z}$, we define
$\Delta_{L}=|a|\prod_{i=1}^{k}|ab_{i}-ba_{i}|.$
We note that if $a_{i}=a$ for all $1\leq i\leq k$, then
$\Delta_{L}=|a|^{k+1}\prod_{i=1}^{k}|b-b_{i}|.$
Modern application of the sieve methods involves the sums
$\sum_{(a,b)\in\Omega}\frac{\Delta_{L}}{\varphi(\Delta_{L})}$
(see, for example, [2]). Here $(a,b)$ denotes a vector and $\Omega$ is a finite set in $\mathbb{Z}^{2}$. We prove
Corollary 1.2.
Let $s$, $a$, $k\geq 2$, $b_{1}<\ldots<b_{k}$ be positive integers. Let $x\geq 3$, $(\ln x)^{-9/10}\leq\eta\leq 1$, and $|b_{i}|\leq\ln x$ for all $1\leq i\leq k$. We set
$f(b)=(b-b_{1})\dots(b-b_{k}).$
Then the following statements hold.
1) If $k\geq\ln\ln x$ or $s\leq k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\ln x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}\eta\ln x.$ (1.5)
2) If $2\leq k<\ln\ln x$ and $s>k$, then
$\sum_{\begin{subarray}{c}|b|\leq\eta\ln x\\ b\,\neq\,b_{1},\ldots,b_{k}\end{subarray}}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}c\,\frac{a}{\varphi(a)}\ln s\Big{)}^{s}\eta\ln x.$ (1.6)
Here $c>0$ is an absolute constant.
Corollary 1.2 extends a result of Maynard ([2, Lemma 8.1]) which showed the same result but with $s=1$ and extends Theorem 1.4 in [3] which showed the inequality (1.6) but with an upper bound
$\Big{(}c\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}s!\,\eta\ln x.$
From Theorem 1.2 we obtain
@@
$\sum_{p\leq x}\Big{(}\frac{p-1}{\varphi(p-1)}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)\pi(x)$ (1.7)
@@
#$\#\big{\{}p\leq x:\frac{p-1}{\varphi(p-1)}>t\big{\}}\leq c_{1}\textup{exp}(-\textup{exp}(c_{2}t))\pi(x).$ (1.8)
@@ 2. Notation
The symbol $b|a$ means that $b$ divides $a$. For fixed $a$ the sum $\sum_{b|a}$ and the product $\prod_{b|a}$ should be interpreted as being over all positive divisors of $a$. If $x$ is a real number, then $[x]$ denotes its integral part and $\{x\}=x-[x]$.
3. Proof of Theorems 1.1 and 1.2
@@
It is easy to see that $f(n)$ is supported on $\mathcal{D}$. Suppose that $m,n\in\mathcal{D}$ and $(m,n)=1$. For any $d_{1},\ldots,d_{s}\in\mathcal{D}$ such that $[d_{1},\ldots,d_{s}]=mn$ there are unique $d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}}\in\mathcal{D},$ $i=1,\ldots,s,$ such that $d_{i}^{{}^{\prime}}d_{i}^{{}^{\prime\prime}}=d_{i}$, $(d_{i}^{{}^{\prime}},d_{i}^{{}^{\prime\prime}})=1$, $[d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n$, and $[d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m$. Hence
$f(nm)=\sum_{\begin{subarray}{c}d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}\in\mathcal{D}:\\ [d_{1}^{{}^{\prime}},\ldots,d_{s}^{{}^{\prime}}]=n\end{subarray}}\frac{\mu^{2}(d_{1}^{{}^{\prime}})\ldots\mu^{2}(d_{s}^{{}^{\prime}})}{d_{1}^{{}^{\prime}}\ldots d_{s}^{{}^{\prime}}}\sum_{\begin{subarray}{c}d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}\in\mathcal{D}:\\ [d_{1}^{{}^{\prime\prime}},\ldots,d_{s}^{{}^{\prime\prime}}]=m\end{subarray}}\frac{\mu^{2}(d_{1}^{{}^{\prime\prime}})\ldots\mu^{2}(d_{s}^{{}^{\prime\prime}})}{d_{1}^{{}^{\prime\prime}}\ldots d_{s}^{{}^{\prime\prime}}}=f(n)f(m).$
Since $f(n)=0$ for any $n\notin\mathcal{D}$, we obtain that $f(n)$ is a multiplicative function. It is clear that
@@
Since $\omega(n)\leq K/g(n)$ for any $n\in\mathcal{D}$, we have
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$ $\leq K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\sum_{n\in\mathcal{D}}\frac{f(n)}{g(n)}=K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{f(p)}{g(p)}\Big{)}$
$=K\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
Theorem 1.1 is proved. ∎
Proof of Theorem 1.2..
Since $g(p)>0$ for any prime $p$, from Theorem 1.1 we obtain
$\sum_{n=1}^{N}\bigg{(}\frac{a_{n}}{\varphi(a_{n})}\bigg{)}^{s}$ $\leq\gamma N\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq\gamma N\Big{(}\frac{c}{\alpha}\Big{)}^{s}\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$ (3.1)
@@
$\Big{(}1+\frac{1}{p}\Big{)}^{s}-1=s(1+\xi)^{s-1}\frac{1}{p}<s\Big{(}1+\frac{1}{s}\Big{)}^{s-1}\frac{1}{p}<\frac{es}{p}.$ (3.2)
@@
$\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\leq\textup{exp}(s\ln\ln(s+2)+Cs)N,$
where $C$ is a positive constant depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$. The inequality (1.1) is proved.
@@
$\textup{exp}(s\ln\ln(s+2)+Cs)N$ $\geq\sum_{n=1}^{N}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}\geq\sum_{n\leq N:\,a_{n}/\varphi(a_{n})>t}\Big{(}\frac{a_{n}}{\varphi(a_{n})}\Big{)}^{s}$
@@
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(s\ln\ln(s+2)+Cs-s\ln t)N.$ (3.5)
@@
$s=[\textup{exp}(te^{-(C+1)})]+1.$
Then
$s+2=\textup{exp}(te^{-(C+1)})+3-\theta,\text{ where }\theta:=\{\textup{exp}(te^{-(C+1)})\},$
and
$\ln(s+2)$ $=te^{-(C+1)}+\ln\Big{(}1+\frac{3-\theta}{\textup{exp}(te^{-(C+1)})}\Big{)}$
$=te^{-(C+1)}+R_{1},\qquad 0<R_{1}\leq\frac{3}{\textup{exp}(te^{-(C+1)})}.$
We get
$\ln\ln(s+2)$ $=\ln t-(C+1)+\ln\Big{(}1+\frac{e^{(C+1)}R_{1}}{t}\Big{)}$
$=\ln t-(C+1)+R_{2},\qquad 0<R_{2}\leq\frac{3e^{C+1}}{t\,\textup{exp}(te^{-(C+1)})}.$
Therefore
$s\ln\ln(s+2)+Cs-s\ln t=-s+sR_{2},$
where
$0<sR_{2}\leq\frac{3e^{C+1}(\textup{exp}(te^{-(C+1)})+1)}{t\,\textup{exp}(te^{-(C+1)})}\leq\frac{6e^{C+1}}{t}.$
We obtain
$s\ln\ln(s+2)+Cs-s\ln t\leq-\textup{exp}(te^{-(C+1)})+\frac{6e^{C+1}}{t}\leq-\textup{exp}(te^{-(C+1)}/2),$ (3.6)
if $t\geq t_{0}$ (here $t_{0}=t_{0}(C)=t_{0}(\gamma,L,c_{0},\alpha)$ is a positive constant depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$).
We set
@@
Then $c_{1}$ and $c_{2}$ are positive constants depending only on $\gamma$, $L$, $c_{0}$, and $\alpha$. From (3.5) and (3.6) we obtain
#$\#\Big{\{}n\leq N:\frac{a_{n}}{\varphi(a_{n})}>t\Big{\}}\leq\textup{exp}(-\textup{exp}(c_{2}t))N,$
if $t>t_{0}$. If $0<t\leq t_{0}$, then
@@
We set $a_{n}=f(n)$. It is clear that $N=[x]$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. We have
$y=(\ln M)^{\alpha}=(d\ln x+\ln\tau)^{1/2}\leq(\ln x)^{3/4},$
@@
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(2(\ln x)^{3/4})\leq\sqrt{x}<[x]=N.$
@@
$\leq(2c)dn^{1-1/d}\,\frac{N}{n}=\frac{\gamma N}{n^{1/d}},$
where $\gamma=2cd$. The function $g(n)=n^{1/d}$ is multiplicative. Also, $g(p)=p^{1/d}\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$) and
@@
Proof of Corollary 1.2..
We can assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. We set
$\Omega=\{b\in\mathbb{Z}:|b|\leq\eta\ln x,b\neq b_{1},\ldots,b_{k}\}.$
Since $\varphi(mn)\geq\varphi(m)\varphi(n)$ for all positive integers $m$ and $n$, and $a^{k+1}/\varphi(a^{k+1})=a/\varphi(a)$, we obtain
$S=\sum_{b\in\Omega}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}=\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}S_{0}.$ (4.1)
Let $b\in\Omega$. Hence, $|b-b_{i}|\leq 2\ln x$ for all $1\leq i\leq k$, and $|f(b)|\leq(2\ln x)^{k}$. Suppose that $k\geq\ln\ln x$. Since $n/\varphi(n)\leq c\ln\ln(n+2)$ for any positive integer $n$, where $c$ is an absolute positive constant, we obtain
$\frac{|f(b)|}{\varphi(|f(b)|)}\leq c(\ln k+\ln\ln(4\ln x))\leq c_{1}\ln k.$
Hence,
$S\leq\Big{(}\frac{a}{\varphi(a)}\,c_{1}\ln k\Big{)}^{s}\#\Omega\leq\Big{(}c_{2}\,\frac{a}{\varphi(a)}\ln k\Big{)}^{s}\eta\ln x,$
where $c_{2}$ is an absolute positive constant.
Suppose that $2\leq k<\ln\ln x$. We take $\alpha=1/4$, $M=(2\ln x)^{k}$. Then
$y=(\ln M)^{\alpha}\leq 2(\ln\ln x)^{1/2}.$
If $n\in\mathcal{D}$, then
$n\leq\prod_{p\leq y}p\leq\textup{exp}(2y)\leq\textup{exp}(4(\ln\ln x)^{1/2})\leq\eta\ln x.$
For $n\in\mathcal{D}$, we have
$\omega(n)=\#\{b\in\Omega:f(b)\equiv 0\text{ (mod $n$)}\}$ $\leq\Big{(}\frac{2\eta\ln x}{n}+1\Big{)}\prod_{p|n}\min(p,k)$
$\leq\frac{3\eta\ln x}{n}\prod_{p|n}\min(p,k).$
We set $g(1)=1,$
$g(p^{\beta})=\begin{cases}p/\min(p,k),&\text{if $\beta=1$;}\\ 1,&\text{if $\beta\geq 2$};\end{cases}$
and $g(p_{1}^{\beta_{1}}\dots p_{r}^{\beta_{r}})=g(p_{1}^{\beta_{1}})\dots g(p_{r}^{\beta_{r}})$, where $p_{1},\ldots,p_{r}$ are pairwise distinct prime numbers. Then $g(n)$ is a multiplicative function and
$\omega(n)\leq\frac{3\eta\ln x}{g(n)}$
for any $n\in\mathcal{D}$. Since $g(p)>0$ for any prime $p$, from Theorem 1.1 we obtain
$\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}\leq c^{s}(\eta\ln x)\prod_{p\leq y}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$
$\leq c^{s}(\eta\ln x)\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}$ (4.2)
(here $c>0$ is an absolute constant). By (4.1) and (4.2) we have
$S\leq\Big{(}c\,\frac{a}{\varphi(a)}\Big{)}^{s}\eta\ln x\prod_{p}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}.$
Suppose that $s\leq k$ (we recall that $2\leq k<\ln\ln x$). If $p>k$, then $p>s$ and (3.2) holds. Since $g(p)=p/k$, we obtain
$\prod_{p>k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>k}\Big{(}1+\frac{eks}{p^{2}}\Big{)}\leq c_{1}^{s},$
where $c_{1}>0$ is an absolute constant. If $p\leq k$, then $g(p)=1$ and we have
$1+\frac{(1+p^{-1})^{s}-1}{g(p)}=(1+p^{-1})^{s}\leq\textup{exp}\Big{(}\frac{s}{p}\Big{)}.$
Hence
$\prod_{p\leq k}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq k}\frac{s}{p}\Big{)}\leq(c_{2}\ln k)^{s},$
where $c_{2}>0$ is an absolute constant. The inequality (1.5) is proved.
Suppose that $s>k$ and $2\leq k<\ln\ln x$. Since $g(p)\geq p/k>p/s$ for any prime $p$, we obtain
$\prod_{p>s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\prod_{p>s}\Big{(}1+\frac{es^{2}}{p^{2}}\Big{)}\leq c_{3}^{s}.$
Since $g(p)\geq 1$ for any prime $p$, we have
$\prod_{p\leq s}\Big{(}1+\frac{(1+p^{-1})^{s}-1}{g(p)}\Big{)}\leq\textup{exp}\Big{(}\sum_{p\leq s}\frac{s}{p}\Big{)}\leq(c_{4}\ln s)^{s}.$
The inequality (1.6) is proved. This completes the proof of Corollary 1.2.
Proof of Corollary 1.3..
We set $a_{n}=p_{n}-1$, where $p_{n}$ is the $n$th prime. It is clear that $N=\pi(x)$. We take $\alpha=1/2$, $M=x$. We can assume that $x\geq x_{0}$, where $x_{0}$ is a large absolute constant. Then $y=(\ln x)^{1/2}$. If $n\in\mathcal{D}$, then
By the Brun - Titchmarsh inequality (see, for example, [4, Exercise 9.3.13]) we have
$\omega(n)=\#\{p\leq x:p\equiv 1\text{ {(}mod $n${)}}\}\leq\frac{3x}{\varphi(n)\ln(2x/n)}\leq\gamma\,\frac{\pi(x)}{\varphi(n)},$
where $\gamma$ is an absolute positive constant. Since $\varphi(n)$ is a multiplicative function, $\varphi(p)=p-1\geq 1$ for any prime $p$ (hence, we can take $c_{0}=1$), and
$L=\sum_{p}\frac{1}{p\varphi(p)}=\sum_{p}\frac{1}{p(p-1)}<\infty,$
we can apply Theorem 1.2. Since $\gamma$, $L$, $c_{0}$, and $\alpha$ are absolute positive constants, the constants $C$, $c_{1}$, and $c_{2}$ in (1.7) and (1.8) are absolute and positive. Corollary 1.3 is proved. ∎